OpenStudy (anonymous):

3x-4 = (2x+4)/3 Show steps please?

OpenStudy (anonymous):

I got 13... but I am not sure that's right. Can any one confirm if I am doing this right?

OpenStudy (anonymous):

3x-4 = (2x+4)/3 3(3x-4) = (2x+4) 9x - 12 = 2x + 4 9x - 16 = 2x 7x - 16 = 0 7x = 16 x = 2.2857

OpenStudy (anonymous):

Oh! I get it! Thank you so much, I almost made a big mistake!

OpenStudy (anonymous):

You're welcome!